There exist a rational number whose product is a whole number. True or false?
1
Expert's answer
2020-09-16T17:58:31-0400

Multiplying two rational numbers produces another rational number.i.e given


"\frac{a} {b}" "×\frac{a} {b}" ="(\frac{a} {b}) ^{2}".


Then all whole numbers are rationals but not all rational numbers are whole numbers, hence there does not exist a rational number whose product is a whole number.